### v2

# Target Speaker Extraction through Comparing Noisy Positive and Negative Audio Enrollments

Abstract
Target speaker extraction focuses on isolating a specific speaker’s voice from an audio mixture containing multiple speakers. To provide information about the target speaker’s identity, prior works have utilized clean audio samples as conditioning inputs. However, such clean audio examples are not always readily available. For instance, obtaining a clean recording of a stranger’s voice at a cocktail party without leaving the noisy environment is generally infeasible. Limited prior research has explored extracting the target speaker’s characteristics from noisy enrollments, which may contain overlapping speech from interfering speakers. In this work, we explore a novel enrollment strategy that encodes target speaker information from the noisy enrollment by comparing segments where the target speaker is talking (Positive Enrollments) with segments where the target speaker is silent (Negative Enrollments). Experiments show the effectiveness of our model architecture, which achieves over 2.1 dB higher SI-SNRi compared to prior works in extracting the monaural speech from the mixture of two speakers. Additionally, the proposed two-stage training strategy accelerates convergence, reducing the number of optimization steps required to reach 3 dB SNR by 60%. Overall, our method achieves state-of-the-art performance in the monaural target speaker extraction conditioned on noisy enrollments.
1 Introduction
In the target speaker extraction task, the model is required to extract the target speaker’s voice from a mixture of multiple speakers and ambient noise. To specify the characteristics of the target speaker, prior works have explored using conditional information of the target speaker in multiple modalities, including visual pan2022selectivelistening ; muse , contextual ma2024clapsep ; typingtolistern , or acoustic modality speakerbeam ; spex ; spex+ ; zhang2024multilevelspeakerrepresentationtarget . Though prior works using conditional information in the acoustic modality have achieved significant good extraction performance, most of these works only considered using clean audio examples of the target speaker he2024hierarchicalspeaker ; meng2024binauralselective ; zhao2024continuoustargetspeech ; pham2024wannahearvoiceadaptive . This strong assumption prevents these models from performing well when only noisy audio examples are available. For example, consider a cocktail party, where the user meets a stranger who has not provided an audio sample before. To extract the target stranger speaker’s voice from the noisy environment to assist in clear conversation, the user will have to ask the speaker to step outside to record the clean audio enrollment. Enrolling target speakers in this way is often impractical in real-world applications.
Figure 1: a) Task scenario explored in the work. Users identify a speaker of interest in an audio mixture, and labeling when the target speaker speaks (Positive Enrollment) or remains silent (Negative Enrollment) in the audio mixture. b) Decomposition of each speaker’s voice in the audio mixture. Due to the stochasticity in human conversation, the interfering speaker will either remain silent in some of the segments in the Positive Enrollment or speak in the Negative Enrollment, leaving the target speaker the only speaker who talks throughout the Positive Enrollments but not in the Negative Enrollment. c) The model performs self-attention between the encoded Positive and Negative Enrollments to extract the target speaker’s characteristic, which serves as the conditional information for the following extraction model. The model then extract the target speaker from the Audio Mixture.
Although prior works have attempted to perform target speaker extraction using noisy enrollments, they either still exploit the timesteps where only the target speaker is speaking in the noisy enrollment adenet ; zhang2024ortse , or assume that the user has participated in the audio mixture recording Chen2024tce ; Veluri2024lookonce , limiting extraction from arbitrary audio mixtures. In this work, we present a method for performing target speaker extraction conditioned on noisy audio enrollments where both the target speaker and the interfering speakers overlap, as shown in Figure 1 (a). To resolve the ambiguity of which speaker in the noisy enrollment is the target speaker, we exploit the stochastic nature of different speakers’ speaking patterns, which allows us to assume that speakers won’t start and stop talking in perfect synchrony in a noisy conversational environment [^1] We quantitatively verify the feasibility of this assumption in Appendix A.. As shown in Figure 1 (b), based on this assumption, we structure the enrollment input as a combination of two segments: a Positive Enrollment (where the target speaker is speaking) and a Negative Enrollment (where the target speaker remains silent). During training, we encourage the model to encode the identity of the speaker who consistently speaks during Positive Enrollments and remains silent during Negative Enrollments. Interfering speakers, from the assumption above, will be inactive during parts of the Positive Enrollment (e.g., Blue speaker in segment C) and/or present during the Negative Enrollment (e.g., Red and Green speakers in segments A and D). By exploiting these temporal misalignments, the model can distinguish and encode the target speaker’s characteristics, despite the presence of interfering speech and noise throughout both enrollments. We detail the problem formulation and the training pipeline to achieve this in Section 3.
By performing target speaker extraction in this manner, our model supports a broad range of real-world applications. Enrollment could easily be achieved in reality by pressing a button on an app or tapping an earbud to indicate a positive segment. Note that precise labeling is not required. Similarly, for offline speaker extraction (e.g., from an audio recording of a meeting), a few seconds of audio can be easily labeled as Positive or Negative samples.
In conclusion, our contributions include:
1. Exploring a novel enrollment strategy for target speaker extraction with noisy positive/negative enrollments and imprecise labels.
2. Design of a two-branch encoder and an associated two-stage training strategy for the task. Ablation experiments show that the proposed training method allows our model to achieve the same level of performance (3.0 dB SNR on the validation set) with 60% fewer optimization steps.
3. Demonstrate and discuss our model’s performance across challenging and realistic application scenarios, including different numbers of interfering speakers, significant overlap between target and interfering speakers in enrollments, and inaccuracies in the positive and negative enrollment labeling.
2 Related Work
2.1 Target Speaker Extraction in Challenging Scenarios
In target speaker extraction, prior works have explored using visual pan2022selectivelistening ; muse , textural ma2024clapsep ; typingtolistern , and audio examples speakerbeam ; spex ; spex+ ; zhang2024multilevelspeakerrepresentationtarget of the target speaker as extraction conditions. Our work belongs to the third category, which learns the acoustic characteristics from given audio examples. Prior works in this category have attempted to improve extraction quality by modifying the fusion method he2024hierarchicalspeaker ; zeng2024useftse , leveraging multi-channel information in the audio mixture input meng2024binauralselective ; pandey2024neurallowlatency , and processing audio in the temporal domain spex ; spex+ .
To improve the models’ robustness and applicability in real-world applications, prior works have explored extracting multiple target speakers’ speech simultaneously rikhye2021multiuservoicefilter ; ma2024clapsep ; zeng2023simultaneousspeech , transferring to extract speech for different languages pham2024wannahearvoiceadaptive , and disentangling irrelevant audio characteristics (e.g., reverberation effect) from the enrollment liu2024tsecurriculum ; Ranjan2018CurriculumLearning ; Heo2024CentroidEstimation ; pandey2023AttentiveTraining ; borsdorf2024wTIMIT2mix ; mu2024selfsuperviseddisentangle . In particular, Zhao et al. zhao2024continuoustargetspeech addressed the problem of variable interfering speaker numbers in the audio mixture to be separated. In our experiment, we also show our model’s generalizability to extract with different numbers of target and interfering speakers. Ranjan et al. Ranjan2018CurriculumLearning adopted curriculum learning to gradually increase the extraction difficulty. After a fixed number of epochs, samples with a lower signal-to-noise ratio (SNR) and higher similarity between target and interfering speakers are added to the training dataset. Our method adopts a similar training paradigm to pretraining the audio encoder. Both ours and the prior work from Ranjan et al. Ranjan2018CurriculumLearning show that the multi-stage training improves model convergence speed and leads to better final performance.
2.2 Target Audio Extraction Conditioned on Enrollments with Interfering Speakers
Though prior works have achieved significant progress in target audio extraction, most of these works assume the availability of clean audio enrollment. However, in real-world application scenarios, such enrollment examples are not necessarily available.
Among those prior works that attempted to extract target speech conditioned on noisy audio enrollments, TCE Chen2024tce explored the turn-taking dynamics in the conversation. TCE model accepts an audio encoding of the user’s clean voice and performs extraction by considering the speakers who cross-talk with the user as the interfering speakers. However, TCE is limited to performing target speaker extraction when the user is participating in the conversation and does not allow specifying audio segments where the target speaker is present. A closely related work to ours is ADEnet adenet , which also utilizes target speaker activity labeling to perform speech extraction. However, ADEnet only validated its performance on two-speaker mixtures with an average overlap ratio of 38.5%, where the enrollment contains mostly clean speech. As a result, it does not explore the challenges of unstable optimization and high overlap ratio between the target speaker and the interferers in the enrollment segment, which we explicitly address in our work. LookOnceToHear Veluri2024lookonce leverages binaural audio to extract the target speaker’s voice. During the enrollment stage, when the user is looking toward the target speaker, the model performs beamforming at a 90-degree azimuth to disambiguate the target speaker from the interferers talking simultaneously.
In comparison to the previous target speaker extraction methods, our model does not assume prior knowledge of any speaker in the mixed audio or the target speaker’s spatial location in the enrollment stage. This allows our model to extract the target speaker from an arbitrary mono-channel sound mixture, and only use easily obtained binary labels of positive and negative audio segments in the temporal domain to extract the target speech. Besides, our model can flexibly perform extraction using either clean or noisy enrollments without retraining and get comparable performance with the previous method trained on clean enrollment, as shown in the Appendix J.
3 Method
3.1 Problem Formulation
In the target speaker extraction (TSE) task, the model first encodes the target speaker’s characteristic from an enrollment input using an encoding branch. Conditioned on the encoding branch output, the model extracts the target speaker’s voice from an Audio Mixture $a^{M}$. To allow the model to extract the target speaker conditioned on noisy enrollments, we formulate our enrollment as a pair of Positive Enrollment $a^{P}$ and Negative Enrollment $a^{N}$. After encoding the target speaker characteristic $F(a^{P},a^{N})$ from the enrollment pair, the model extracts the target speaker’s voice from the audio mixture $G(a^{M},F(a^{P},a^{N}))$.
To simulate the audio mixture in a noisy environment, we model the signal as the sum of an individual speaker’s voice and the background noise. Thus, the noisy Positive Enrollment $a^{P}$, Negative Enrollment $a^{N}$, and the Audio Mixture $a^{M}$ are constructed as:
$a^{M}=\sum_{i\in S^{IM}\cup\{t\}}a^{M}_{i}+n^{M},\,a^{P}=\sum_{i\in S^{IE}\cup\{t\}}a^{P}_{i}+n^{P},\,a^{N}=\sum_{i\in S^{IE}}a^{N}_{i}+n^{N},$ (1)
where $a^{\{P,N,M\}}_{i}$ represents the speech signal of speaker $i$ in the Positive, Negative, and Audio Mixture. $n^{\{P,N,M\}}$ represents the background noise in the three mixtures. $t$ represents the target speaker ID, and $S^{IE},S^{IM}$ represent the interfering speakers in the enrollments and the Audio Mixture, respectively.
3.2 Speaker Disambiguation via Positive and Negative Enrollments
To show how the Positive and Negative Enrollment pair resolves the ambiguity of which speaker is the target speaker in the noisy enrollments, we introduce the following assumption about speakers’ temporal dynamics:
Assumption: In a long audio mixture where multiple speakers speak independently, no two speakers will consistently start and stop talking at exactly the same times throughout the mixture.
This assumption is easily satisfied: For two speakers to always talk simultaneously, one speaker would have to deliberately start and stop speaking at the exact same times as the other, effectively aiming to sabotage the other’s conversation. Such speakers interfering with intention are unlikely to exist in typical cocktail-party scenario, when interfering speakers in the environment are talking independently from the target speaker. We quantitatively verify this assumption on the real-world audio dataset in Appendix A.
Based on the assumption above, the Positive and Negative Enrollments divide interfering speakers into 4 categories, depending on their existence in both enrollments:
1) Negative Interferer (NI): These are speakers who talk consistently throughout the Positive Enrollment and are present in the Negative Enrollment. The model should learn their voice characteristics from the Negative Enrollment and exclude them from the encoding.
2) Positive Interferer (PI): These are speakers who are not present in the Negative Enrollment and talk in a fraction of the Positive Enrollment. The model should identify those segments when PIs are absent in the Positive Enrollment and exclude these speakers in the encoding. We explore how sensitive our model is to this fractional overlap in Section 4.4.
3) Hybrid Interferer (HI): These speakers appear partially in both the Positive and Negative Enrollments. They resemble a combination of Positive and Negative Interferers, providing the model with two potential cues for suppression: their absence in some Positive Enrollment segments and their presence in some of the Negative Enrollment segments.
4) Neglect-Required Interferer (NRI): These speakers appear exclusively in the Negative Enrollment. Similar to the Negative Interferers, the model encoder should avoid include these speakers’ voice characteristic in the extracted embedding.
Since Positive and Negative Interferers represent the two primary scenarios in which the model must learn to distinguish target speakers from interfering ones, we focus our training and evaluation on these two types. We demonstrate our model’s generalizability when Hybrid and Neglect-Required Interferers exist in Appendix H and I, respectively.
3.3 Model Architecture
Figure 2: Encoding and Extraction Branch model architecture and training pipeline.
As shown in Figure 2, we adopt TF-GridNet tfgridnet as the backbone for both our Encoding and Extraction Branches. The TF-GridNet architecture consists of a 2D convolution layer followed by stacks of three TF-GridNet blocks. Each block consists of two BiLSTM modules that capture the inter-frequency and temporal information, followed by a Full-band Self-attention Module designed to capture the long-range frame information between frames. To adapt the original TF-GridNet, which is designed to perform sound separation, to the Target Speaker Extraction task, we introduce an attention-based Encoder Fusion Module and an Extraction Fusion Module. The Encoder Fusion Module performs comparison between Positive and Negative Enrollment embeddings to generate the embedding of the target speaker, while the Extraction Fusion Module integrates enrollment encoding into the extraction model. The model architecture and training method for encoding the target speaker’s feature are shown in Figure 2.
The Positive and Negative Enrollments are first encoded by a pair of TF-GridNet encoders in the Siamese Encoder. Since The goal of both branches in the Siamese Encoder is to encode multiple speakers’ characteristics from the noisy Positive or Negative Enrollments, we share the parameters between the two TF-GridNet encoders of the Siamese Encoder to reduce model parameter size. The Siamese encoder results in two sequences of embeddings $E_{pos}$ and $E_{neg}$ with shapes $[T_{pos},D]$ and $[T_{neg},D]$, respectively.
The Encoder Fusion Module then extracts the target speaker’s voice embedding from the Siamese Encoders’ output $E_{pos}$ and $E_{neg}$. To encode the target speaker’s characteristic that exists throughout the Positive Enrollment but is absent in the Negative Enrollment, we need to perform a pairwise comparison between frame-level embeddings of both enrollments. First, we element-wise add each frame embedding in both sequences with learnable segmentation embeddings, to indicate whether it originates from the Positive or Negative Enrollment. We then adopt a self-attention mechanism, which allows the model to dynamically compute similarity scores between all frame embeddings. These attention scores serve as a soft alignment, enabling each positive frame to attend to other positive frame embeddings to remove the Positive Interferers’ voice characteristic from its encoding, and attend to the negative frame embeddings to remove Negative Interferers’ voice characteristics. After two Full-band Self-attention Modules, the model is able to selectively encode only target-speaker-specific features while suppressing interfering signals from background noise or other speakers. We retain only the embeddings corresponding to the positive enrollment as the extracted target speaker embedding, which reduces the output to a $[T_{\text{pos}},D]$ shaped embedding $E_{\text{fused}}$. This reduction in embedding size ensures the embedding feature is the same size as the pretrained encoder’s output, allowing the model to benefit from the fast convergence brought by training the encoder using knowledge distillation (as shown in Section 4.3). To further reduce the computation load in the extraction branch, we apply non-overlapping average pooling with 40 kernel size along the temporal dimension to $E_{\text{fused}}$, resulting in a $[T_{\text{pos}}//40,D]$ shaped feature $E_{\text{fused-pooled}}$. We discuss the effect of different kernel sizes on the model performance and inference time in Appendix D.
Conditioned on the pooled enrollment encoder output $E_{\text{fused-pooled}}$, we use a TF-GridNet-based Extraction Branch to extract the target speaker’s voice from the Audio Mixture. To perform causal target speaker extraction, we modify the TF-GridNet block by adopting the same modification as LookOnceToHear Veluri2024lookonce . Three causal TF-GridNet blocks are used in the Extraction Branch. Two cross-attention-based Extraction Fusion Blocks are added after the first two causal TF-GridNet blocks to integrate the pooled target speaker embedding $E_{\text{fused-pooled}}$ into the extraction model. In each fusion module, the pooled target speaker embeddings $E_{\text{fused-pooled}}$ serve as the Key and Value, while the output from the previous TF-GridNet block serves as the Query for the fusion module. The output from the last TF-GridNet block passes through a transposed convolution layer and an ISTFT module to generate the speech waveform.
3.4 Training Method
Due to the additional speaker and background noise in the enrollments, the noisy Positive and Negative Enrollment pairs introduce significant variability in the training data in comparison to the clean enrollments. As shown in Section 4.3, when the Encoding and the Extraction branches are trained together using these noisy training samples, the model converges much more slowly and reaches worse final performance. Motivated by the prior works Veluri2024lookonce ; splitrole ; voicefilter that successfully trained extraction models conditioned on the clean target speaker embeddings, we separately train the Encoding and Extraction branches in a two-stage training strategy.
As shown in Figure 2, the first stage trains the Siamese Encoder and the Encoder Fusion Module from scratch, using knowledge distillation, to extract the target speaker’s embedding from the noisy positive and negative audio enrollments. We obtain the ground-truth target speaker’s embedding by passing the clean target speaker’s voice through a trained frozen TF-GridNet encoder Veluri2024lookonce . Formally, the first stage training loss is written as
$L_{\text{stage 1}}=\|E_{\text{clean}}-E_{\text{fused}}\|^{2}\,,$ (2)
where $E_{\text{clean}}$ is the embedding of the clean target speaker voice $a_{\text{clean}}$ encoded by a trained TF-GridNet encoder from Veluri2024lookonce .
In the second stage, we train the Extraction Branch to extract the target speaker from the Audio Mixture. The training loss for the second stage is the negative SNR value of the extracted audio $\hat{a}_{tgt}$ and the ground-truth target speaker speech $a_{tgt}$:
$L_{\text{stage 2}}=-\text{SNR}(\hat{a}_{tgt},a_{tgt}).$ (3)
More training and model details are discussed in Appendix C.
4 Experiment
4.1 Datasets and Baselines
Datasets
We construct our Positive, Negative, and Mixed Audios from the samples in the LibriSpeech dataset librispeech , with background noise $n^{\{M,P,N\}}$ from the WHAM! dataset wham . Following the experiment configuration in LookOnceToHear Veluri2024lookonce , we also construct binaural Positive / Negative / Mixed Audio samples by convolving the speech of each speaker with the binaural RIR data from 4 different datasets: CIPIC cipic , RRBRIR RRBRIR , ASH-Listening-Set ASH-brir , and CATTRIR CATTRIR .
In evaluation, 5000 samples are generated from the LibriSpeech dataset test-clean component using the same strategy as the training samples. We report the average improvement in SNR (SNRi) and SI-SNR (SI-SNRi), and the extracted audio’s STOI and DNSMOS, along with the standard deviation across evaluation samples calculated using the built-in Pytorch function under the assumption that model performance metrics follow a normal distribution. We detail the training and testing data generation in Appendix B.
Baselines
We compare our monaural model’s performance with TCE Chen2024tce , SpeakerBeam speakerbeam , and non-negative matrix factorization (NMF) method NMF . We compare our binaural model’s reverberant target speech extraction performance with the LookOnceToHear model Veluri2024lookonce .
1. TCE Chen2024tce considers speakers whose voices do not overlap with a given speaker as the target speakers and extracts their voices. The model takes in a d-vector embedding of the given speaker and removes all interfering speakers who talk at the same time as the given speaker.
2. LookOnceToHear Veluri2024lookonce is a binaural target speech extraction method conditioned on noisy audio enrollments. This method uses beamforming to extract the target speaker’s characteristics from the noisy Positive Enrollment, which contains audio from multiple additional speakers.
3. SpeakerBeam speakerbeam is a target speech extraction model using clean target speaker enrollment as the extraction condition. We include this model as a baseline to show the influence of having interfering speakers in the audio enrollment on the model performance.
4. NMF NMF is a non-deep-learning-based audio separation technique. Since the method could not perform target speaker extraction conditioned on noisy audio examples, we report its extraction quality by selecting the audio with the highest SNR among its output as its extracted audio.
4.2 Result on Monaural and Binaural Reverberant Target Speech Extraction
Table 1: Comparison between our method and the baselines’ performance when extracting from an audio mixture of 2-3 speakers conditioned on an enrollment mixture of 2-4 speakers. NMF’s performance does not change as the number of speakers in the Enrollments changes, as it performs unconditional sound separation. Differences in speaker similarity across combinations substantially impact extraction difficulty, leading to the high standard deviation across samples.
Method Audio Type Metric 2 Speakers in Extraction 3 Speakers in Extraction
2 Speakers 3 Speakers 4 Speakers 2 Speakers 3 Speakers 4 Speakers
in Enroll in Enroll in Enroll in Enroll in Enroll in Enroll
NMF NMF Mono SNRi 4.24$\pm$1.60 5.65$\pm$1.85
SI-SNRi -1.65$\pm$3.78 -1.50$\pm$3.57
STOI 0.362$\pm$0.139 0.296$\pm$0.114
DNSMOS 1.56$\pm$0.44 1.55$\pm$0.41
SpeakerBeam speakerbeam Mono SNRi 4.78$\pm$2.87 4.84$\pm$2.83 4.84$\pm$2.85 5.91$\pm$2.52 5.58$\pm$2.48 5.72$\pm$2.48
SI-SNRi 2.49$\pm$12.19 1.75$\pm$11.89 0.51$\pm$11.86 1.11$\pm$8.46 -0.33$\pm$8.70 -1.07$\pm$8.84
STOI 0.617$\pm$0.271 0.595$\pm$0.270 0.564$\pm$0.268 0.486$\pm$0.212 0.451$\pm$0.212 0.440$\pm$0.213
DNSMOS 1.72$\pm$0.57 1.73$\pm$0.59 1.71$\pm$0.61 1.57$\pm$0.51 1.51$\pm$0.40 1.49$\pm$0.52
TCE Chen2024tce Mono SNRi 8.48$\pm$2.34 7.84$\pm$2.56 7.52$\pm$2.68 8.47$\pm$2.37 8.57$\pm$2.34 8.07$\pm$2.40
SI-SNRi 6.67$\pm$3.69 5.57$\pm$4.47 4.77$\pm$5.34 6.63$\pm$3.74 5.13$\pm$4.07 4.10$\pm$4.60
STOI 0.682$\pm$0.120 0.650$\pm$0.140 0.624$\pm$0.160 0.682$\pm$0.120 0.551$\pm$0.143 0.521$\pm$0.151
DNSMOS 1.85$\pm$0.44 1.84$\pm$0.44 1.83$\pm$0.44 1.85$\pm$0.44 1.65$\pm$0.41 1.65$\pm$0.41
Mono SNRi 8.53$\pm$2.30 8.39$\pm$2.37 8.37$\pm$2.50 9.03$\pm$2.23 9.00$\pm$2.31 8.89$\pm$2.38
Ours SI-SNRi 6.70$\pm$3.43 6.45$\pm$3.74 6.31$\pm$3.84 6.25$\pm$3.35 6.06$\pm$3.68 5.76$\pm$3.95
(Film fusion) STOI 0.681$\pm$0.111 0.675$\pm$0.123 0.665$\pm$0.127 0.581$\pm$0.123 0.578$\pm$0.129 0.565$\pm$0.133
DNSMOS 1.92$\pm$0.39 1.90$\pm$0.39 1.90$\pm$0.41 1.69$\pm$0.38 1.68$\pm$0.38 1.67$\pm$0.38
Mono SNRi 10.14$\pm$2.57 9.93$\pm$2.71 9.79$\pm$2.86 10.42$\pm$2.49 10.37$\pm$2.64 10.22$\pm$2.79
Ours SI-SNRi 8.85$\pm$3.67 8.54$\pm$4.01 8.30$\pm$4.47 8.42$\pm$3.62 8.29$\pm$4.06 7.82$\pm$4.62
(Monaural) STOI 0.758$\pm$0.107 0.749$\pm$0.121 0.742$\pm$0.126 0.668$\pm$0.123 0.665$\pm$0.130 0.648$\pm$0.146
DNSMOS 2.14$\pm$0.37 2.13$\pm$0.37 2.02$\pm$0.38 1.93$\pm$0.37 1.92$\pm$0.38 1.90$\pm$0.38
Bi SNRi 9.36$\pm$3.56 8.90$\pm$3.57 9.12$\pm$3.57 9.30$\pm$3.39 9.28$\pm$3.49 9.30$\pm$3.56
LookOnceToHear SI-SNRi 7.75$\pm$5.35 7.24$\pm$5.43 7.53$\pm$5.35 6.58$\pm$5.37 6.49$\pm$5.62 6.51$\pm$5.64
Veluri2024lookonce STOI 0.736$\pm$0.152 0.723$\pm$0.162 0.732$\pm$0.153 0.620$\pm$0.175 0.610$\pm$0.180 0.611$\pm$0.187
DNSMOS 1.79$\pm$0.56 1.77$\pm$0.55 1.76$\pm$0.55 1.59$\pm$0.47 1.60$\pm$0.48 1.58$\pm$0.49
Bi SNRi 9.84$\pm$3.57 9.60$\pm$3.57 9.59$\pm$3.63 9.81$\pm$3.28 9.78$\pm$3.23 9.83$\pm$3.30
Ours SI-SNRi 8.18$\pm$4.81 7.84$\pm$5.20 7.75$\pm$5.17 6.76$\pm$5.27 6.72$\pm$5.08 6.71$\pm$5.50
(Binaural) STOI 0.735$\pm$0.150 0.729$\pm$0.155 0.720$\pm$0.159 0.608$\pm$0.177 0.605$\pm$0.177 0.605$\pm$0.182
DNSMOS 1.82$\pm$0.60 1.80$\pm$0.59 1.80$\pm$0.59 1.59$\pm$0.50 1.60$\pm$0.49 1.59$\pm$0.48
We compare our model with baseline methods under scenarios where different numbers of speakers are present. Table 1 shows the monaural target speaker extraction performance. SpeakerBeam speakerbeam shows significantly worse performance, especially as the number of interfering speaker increase in the enrollment. TCE Chen2024tce achieves better performance than SpeakerBeam speakerbeam baseline since it does not rely on the clean target speaker’s enrollment. However, apart from the STOI metric in one scenario, the TCE model shows worse performance than our method under all different numbers of interfering speakers in the Audio Mixture and in enrollments.
In the multi-channel target speech extraction, prior work Veluri2024lookonce has explored using beamforming to extract the target speaker’s characteristics. In this experiment, we show that additionally including the Negative Enrollment helps improve model performance. As shown in Table 1, in comparison to the LookOnceToHear Veluri2024lookonce baseline, based on two-sample t-tests over the 5000 test samples at the 90% confidence level, our model achieves statistically significant improvements in SNRi and SI-SNRi under all conditions. The worse STOI performance might be caused by the difference in parameter size, as shown in Appendix C. In contrast to the monaural experiment results, the binaural models show lower performance in most of the metrics and scenarios. This performance drop might be caused by the difference in training objective, since the binaural models need to predict the reverberant binaural audio, while the monaural models are trained to extract non-reverberant speech (to achieve fair comparison with the prior works focusing on non-reverberant extraction). The DNSMOS metric is also affected by the reverberation effect in the prediction, thus providing less effective information for comparing model performance. We include it here only for completeness.
4.3 Ablation Study
Effectiveness of the two-stage model training
Figure 3: Validation loss values for the optimization step. The curve for the two-stage training begins at the 200k step to account for the 200k optimization steps performed during the first training stage.
In this section, we show the effectiveness of the two-stage training by showing the validation loss curve of our model with and without the first training stage. As shown in Figure 3, end-to-end training reaches 3 dB SNR on the validation set after 600k optimization steps (around 125 hours), while the combined training time for the two-staged training takes 240k optimization steps (around 50 hours) based on a single Nvidia A10 24GB GPU.
Such a performance difference might be caused by the high difficulty in encoding the expected target speaker from the noisy enrollment pair. Since its very challenging for the model encoder to encode the target speaker from the noisy enrollment pair, it is difficult for the model to draw similarity between its extraction ground-truth and the noisy pair of enrollment input, resulting in the end-to-end learning providing little guidance to the encoder on which speaker is expected to be encoded. In contrast, the two-stage training explicitly guides the model to extract the target speaker’s embedding in the first stage, significantly accelerating convergence.
Cross-Attention based Fusion over Film Fusion Method
Film fusion is widely applied in the prior target audio extraction works film-fusion ; Chen2024tce ; Veluri2024lookonce . However, Film fusion passes the condition embedding through a linear layer, and element-wise multiplies the output with the input tensor to perform fusion. This limits the model to use a fixed embedding size, which might not be capable of encoding the fine-grained details of the target speaker’s voice characteristics. In this experiment, we show that our attention-based fusion method is more optimal for our target speech extraction task.
We keep the rest of the model architecture intact and only change the model fusion method. In particular, we perform global average pooling on the encoder head output along the temporal dimension to obtain an embedding as the condition input to the Film fusion block. As shown in Table 1, our attention-based fusion method achieves higher performance in all application scenarios. This shows that the cross-attention-based fusion method is more preferable for the target speech extraction conditioned on noisy enrollments.
4.4 Model Performance in Challenging Application Scenarios
Model Performance under different PI and NI’s speech length
As the Positive Interferer (PI)’s speech duration in the Positive Enrollment increases, fewer frames are available for the model to distinguish the Target Speaker from the PI. Similarly, reducing the Negative Interferer (NI)’s speech length in the Negative Enrollment results in less available conditional information to remove NI’s voice characteristics in the target speaker encoding. In this section, we explicitly examine how the model performance varies with the duration of the PI and NI’s speech in the enrollment.
We focus on the scenario where one PI and one NI exist in the enrollments, and the Audio Mixture contains the same interferer as those in the enrollments. Let the length of the PI’s speech in the Positive Enrollment be $l_{pos}$, and the NI’s speech in the Negative Enrollment be $l_{neg}$. In Figure 4, we report the average model performance when extracting 2000 test samples with $l_{neg}$ randomly sampled between 1s to 3s and $l_{pos}$ takes different values between 0s to 3s. We evaluate the extracted audio’s SI-SNRi with respect to the target speaker and the PI’s speech in the Audio Mixture. When $l_{pos}=3$, the model fails to distinguish the two speakers as expected, since both speakers have the same speech length in each enrollment. As $l_{pos}$ decreases below 2.7s, the model correctly identifies the target speaker’s characteristic, as shown by the increasing SI-SNRi of the extracted audio w.r.t. the target speaker’s speech and the decreasing SI-SNRi w.r.t. the PI’s speech. Alternatively, we randomly sample $l_{pos}$ between 1s to 3s and evaluate the model performance when $l_{neg}$ takes a value between 0s to 3s. As shown in Figure 5, the model correctly extracts the target speaker and removes the NI’s speech when $l_{neg}$ increases above 0.3s. These experiments demonstrate that the model indeed leverages the differences in speakers’ speech length in both enrollments to encode the target speaker, and could distinguish speakers from as short as 0.3 seconds of misalignment between their speech.
Figure 4: Model performance w.r.t. length of Positive Interferer in the Positive Enrollment.
Figure 5: Model performance w.r.t. length of Negative Interferer in the Negative Enrollment.
Model performance under inaccurate Positive and Negative Labeling
Figure 6: Box plot of model performance under inaccurate Positive and Negative Enrollment labeling from users. Samples 1-5 contain 2 speakers in enrollments and Audio Mixture, and 6-10 contain 3 speakers. The black dots represent the model performance when using the ground truth Positive and Negative timestep labelings, and the red squares represent outliers.
In the real-world application scenario, inaccurate Positive and Negative Enrollment labeling from the users could lead to unwanted inclusion of the target speaker’s voice in the Negative Enrollments, or including timesteps when the target speaker remains silent in the Positive Enrollment. Both cases violate our model input assumption. In this section, we demonstrate our model’s tolerance to the inaccuracy in the user-provided Enrollments’ labeling.
We randomly select 2-3 speakers and overlay their voices so that each speaker spans 5-7 seconds of a 10-second audio mixture. For 10 such randomly generated samples, we instruct 10 users to label the start and stop timesteps of a certain speaker following a speaker hint, e.g, the same speaker heard in another clean audio example. The segments not labeled by the user are considered negative enrollments. Using the enrollment labeling provided by users, we perform extraction on another audio mixture containing the target speaker. We report the SI-SNRi of each sample across different users’ labeling, as well as the SI-SNRi if the ground truth labeling is used. As shown in Figure 6, our model shows little performance degradation when using the inaccurate positive enrollment labeling provided by the user, despite each start and stop timestep of the positive enrollment being on average $0.30\pm 0.19$ seconds away from the ground truth timestep. In addition, the model shows a small variation in metric value across different users’ labeling. This validates our model’s robustness to the inaccuracy in different users’ labeling.
5 Conclusion and Limitation
In this paper, we present a method for performing target speech extraction conditioned on noisy enrollments where interfering speakers’ speech overlaps significantly with the target speaker. We optimize our model to encode the target speaker from easily obtained noisy positive and negative enrollments, without using any clean audio enrollment input. Experiments show that our method achieves SOTA performance under the challenging and realistic application scenario.
Though the proposed two-stage training strategy for the encoding branch significantly accelerates convergence, it may limit the encoder’s performance to that of the frozen encoder used for knowledge distillation. Further work on encoding branch training is required to remove such potential performance limitations while maintaining the high convergence speed. In addition, our model extraction still includes artifacts noticeable to the human ear, as shown by the low DNSMOS score reported. This hinders the model’s applicability to downstream tasks such as audio recognition. Further work is required to reduce the artifacts, for example, by employing more advanced extraction branch architectures. Finally, although our model can flexibly perform extraction using clean or noisy audio enrollments, it does not outperform SOTA models explicitly trained to extract conditioned on clean enrollments. Further work is required to develop models with strong generalizability across both enrollment strategies.