### v3

# Target Speaker Extraction through Comparing Noisy Positive and Negative Audio Enrollments

Abstract
Target speaker extraction focuses on isolating a specific speaker’s voice from an audio mixture containing multiple speakers. To provide information about the target speaker’s identity, prior works have utilized clean audio samples as conditioning inputs. However, such clean audio examples are not always readily available. For instance, obtaining a clean recording of a stranger’s voice at a cocktail party without leaving the noisy environment is generally infeasible. Limited prior research has explored extracting the target speaker’s characteristics from noisy enrollments, which may contain overlapping speech from interfering speakers. In this work, we explore a novel enrollment strategy that encodes target speaker information from the noisy enrollment by comparing segments where the target speaker is talking (Positive Enrollments) with segments where the target speaker is silent (Negative Enrollments). Experiments show the effectiveness of our model architecture, which achieves over 2.1 dB higher SI-SNRi compared to prior works in extracting the monaural speech from the mixture of two speakers. Additionally, the proposed two-stage training strategy accelerates convergence, reducing the number of optimization steps required to reach 3 dB SNR by 60%. Overall, our method achieves state-of-the-art performance in the monaural target speaker extraction conditioned on noisy enrollments. Our implementation is available at https://github.com/xu-shitong/TSE-through-Positive-Negative-Enroll .
1 Introduction
In the target speaker extraction task, the model is required to extract the target speaker’s voice from a mixture of multiple speakers and ambient noise. To specify the characteristics of the target speaker, prior works have explored using conditional information of the target speaker in multiple modalities, including visual pan2022selectivelistening; muse, contextual ma2024clapsep; typingtolistern, or acoustic modality speakerbeam; spex; spex+; zhang2024multilevelspeakerrepresentationtarget. Though prior works using conditional information in the acoustic modality have achieved significant good extraction performance, most of these works only considered using clean audio examples of the target speaker he2024hierarchicalspeaker; meng2024binauralselective; zhao2024continuoustargetspeech; pham2024wannahearvoiceadaptive. This strong assumption prevents these models from performing well when only noisy audio examples are available. For example, consider a cocktail party, where the user meets a stranger who has not provided an audio sample before. To extract the target stranger speaker’s voice from the noisy environment to assist clear conversation, the user will have to ask the speaker to step outside to record the clean audio enrollment. Enrolling target speakers in this way is often impractical in real-world applications.
Figure 1: a) Task scenario explored in the work. Users identify a speaker of interest in an audio mixture, and labeling when the target speaker speaks (Positive Enrollment) or remains silent (Negative Enrollment) in the audio mixture. b) Decomposition of each speaker’s voice in the audio mixture. Due to the stochasticity in human conversation, the interfering speaker will either remain silent in some of the segments in the Positive Enrollment or speak in the Negative Enrollment, leaving the target speaker the only speaker who talks throughout the Positive Enrollments but not in the Negative Enrollment. c) The model performs self-attention between the encoded Positive and Negative Enrollments to extract the target speaker’s characteristic, which serves as the conditional information for the following extraction model. The model then extract the target speaker from the Audio Mixture.
Although prior works have attempted to perform target speaker extraction using noisy enrollments, they either still exploit the timesteps where only the target speaker is speaking in the noisy enrollment adenet; zhang2024ortse, or assume that the user has participated in the audio mixture recording Chen2024tce; Veluri2024lookonce, limiting extraction from arbitrary audio mixtures. In this work, we present a method for performing target speaker extraction conditioned on noisy audio enrollments where both the target speaker and the interfering speakers overlap, as shown in Figure 1 (a). To resolve the ambiguity of which speaker in the noisy enrollment is the target speaker, we exploit the stochastic nature of different speakers’ speaking patterns, which allows us to assume that speakers won’t start and stop talking in perfect synchrony in a noisy conversational environment [^1] We quantitatively verify the feasibility of this assumption in Appendix A.. As shown in Figure 1 (b), based on this assumption, we structure the enrollment input as a combination of two segments: a Positive Enrollment (where the target speaker is speaking) and a Negative Enrollment (where the target speaker remains silent). During training, we encourage the model to encode the identity of the speaker who consistently speaks during Positive Enrollments and remains silent during Negative Enrollments. Interfering speakers, from the assumption above, will be inactive during parts of the Positive Enrollment (e.g., Blue speaker in segment C) and/or present during the Negative Enrollment (e.g., Red and Green speakers in segments A and D). By exploiting these temporal misalignments, the model can distinguish and encode the target speaker’s characteristics, despite the presence of interfering speech and noise throughout both enrollments. We detail the problem formulation and the training pipeline to achieve this in Section 3.
By performing target speaker extraction in this manner, our model supports a broad range of real-world applications. Enrollment could easily be achieved in reality by pressing a button on an app or tapping an earbud to indicate a positive segment. Note that precise labeling is not required. Similarly, for offline speaker extraction (e.g., from an audio recording of a meeting), a few seconds of audio can be easily labeled as Positive or Negative samples.
In conclusion, our contributions include:
1. Exploring a novel enrollment strategy for target speaker extraction with noisy positive/negative enrollments and imprecise labels.
2. Design of a two-branch encoder and an associated two-stage training strategy for the task. Ablation experiments show that the proposed training method allows our model to achieve the same level of performance (3.0 dB SNR on the validation set) with 60% fewer optimization steps.
3. Demonstrate and discuss our model’s performance across challenging and realistic application scenarios, including different numbers of interfering speakers, significant overlap between target and interfering speakers in enrollments, and inaccuracies in the positive and negative enrollment labeling.
2 Related Work
2.1 Target Speaker Extraction in Challenging Scenarios
In target speaker extraction, prior works have explored using visual pan2022selectivelistening; muse, textural ma2024clapsep; typingtolistern, and audio examples speakerbeam; spex; spex+; zhang2024multilevelspeakerrepresentationtarget of the target speaker as extraction conditions. Our work belongs to the third category, which learns the acoustic characteristics from given audio examples. Prior works in this category have attempted to improve extraction quality by modifying the fusion method he2024hierarchicalspeaker; zeng2024useftse, leveraging multi-channel information in the audio mixture input meng2024binauralselective; pandey2024neurallowlatency, and processing audio in the temporal domain spex; spex+.
To improve the models’ robustness and applicability in real-world applications, prior works have explored extracting multiple target speakers’ speech simultaneously rikhye2021multiuservoicefilter; ma2024clapsep; zeng2023simultaneousspeech, transferring to extract speech for different languages pham2024wannahearvoiceadaptive, and disentangling irrelevant audio characteristics (e.g., reverberation effect) from the enrollment liu2024tsecurriculum; Ranjan2018CurriculumLearning; Heo2024CentroidEstimation; pandey2023AttentiveTraining; borsdorf2024wTIMIT2mix; mu2024selfsuperviseddisentangle. In particular, Zhao et al. zhao2024continuoustargetspeech addressed the problem of variable interfering speaker numbers in the audio mixture to be separated. In our experiment, we also show our model’s generalizability to extract with different numbers of target and interfering speakers. Ranjan et al. Ranjan2018CurriculumLearning adopted curriculum learning to gradually increase the extraction difficulty. After a fixed number of epochs, samples with a lower signal-to-noise ratio (SNR) and higher similarity between target and interfering speakers are added to the training dataset. Our method adopts a similar training paradigm to pretraining the audio encoder. Both ours and the prior work from Ranjan et al. Ranjan2018CurriculumLearning show that the multi-stage training improves model convergence speed and leads to better final performance.
2.2 Target Audio Extraction Conditioned on Enrollments with Interfering Speakers
Though prior works have achieved significant progress in target audio extraction, most of these works assume the availability of clean audio enrollment. However, in real-world application scenarios, such enrollment examples are not necessarily available.
Among those prior works that attempted to extract target speech conditioned on noisy audio enrollments, TCE Chen2024tce explored the turn-taking dynamics in the conversation. TCE model accepts an audio encoding of the user’s clean voice and performs extraction by considering the speakers who cross-talk with the user as the interfering speakers. However, TCE is limited to performing target speaker extraction when the user is participating in the conversation and does not allow specifying audio segments where the target speaker is present. A closely related work to ours is ADEnet adenet, which also utilizes target speaker activity labeling to perform speech extraction. However, ADEnet only validated its performance on two-speaker mixtures with an average overlap ratio of 38.5%, where the enrollment contains mostly clean speech. As a result, it does not explore the challenges of unstable optimization and high overlap ratio between the target speaker and the interferers in the enrollment segment, which we explicitly address in our work. LookOnceToHear Veluri2024lookonce leverages binaural audio to extract the target speaker’s voice. During the enrollment stage, when the user is looking toward the target speaker, the model performs beamforming at a 90-degree azimuth to disambiguate the target speaker from the interferers talking simultaneously.
In comparison to the previous target speaker extraction methods, our model does not assume prior knowledge of any speaker in the mixed audio or the target speaker’s spatial location in the enrollment stage. This allows our model to extract the target speaker from an arbitrary mono-channel sound mixture, and only use easily obtained binary labels of positive and negative audio segments in the temporal domain to extract the target speech. Besides, our model can flexibly perform extraction using either clean or noisy enrollments without retraining and get comparable performance with the previous method trained on clean enrollment, as shown in the Appendix J.
3 Method
3.1 Problem Formulation
In the target speaker extraction (TSE) task, the model first encodes the target speaker’s characteristic from an enrollment input using an encoding branch. Conditioned on the encoding branch output, the model extracts the target speaker’s voice from an Audio Mixture $a^{M}$. To allow the model to extract the target speaker conditioned on noisy enrollments, we formulate our enrollment as a pair of Positive Enrollment $a^{P}$ and Negative Enrollment $a^{N}$. After encoding the target speaker characteristic $F(a^{P},a^{N})$ from the enrollment pair, the model extracts the target speaker’s voice from the audio mixture $G(a^{M},F(a^{P},a^{N}))$.
To simulate the audio mixture in a noisy environment, we model the signal as the sum of an individual speaker’s voice and the background noise. Thus, the noisy Positive Enrollment $a^{P}$, Negative Enrollment $a^{N}$, and the Audio Mixture $a^{M}$ are constructed as:
$a^{M}=\sum_{i\in S^{IM}\cup\{t\}}a^{M}_{i}+n^{M},\,a^{P}=\sum_{i\in S^{IE}\cup\{t\}}a^{P}_{i}+n^{P},\,a^{N}=\sum_{i\in S^{IE}}a^{N}_{i}+n^{N},$ (1)
where $a^{\{P,N,M\}}_{i}$ represents the speech signal of speaker $i$ in the Positive, Negative, and Audio Mixture. $n^{\{P,N,M\}}$ represents the background noise in the three mixtures. $t$ represents the target speaker ID, and $S^{IE},S^{IM}$ represent the interfering speakers in the enrollments and the Audio Mixture, respectively.
3.2 Speaker Disambiguation via Positive and Negative Enrollments
To show how the Positive and Negative Enrollment pair resolves the ambiguity of which speaker is the target speaker in the noisy enrollments, we introduce the following assumption about speakers’ temporal dynamics:
Assumption: In a long audio mixture where multiple speakers speak independently, no two speakers will consistently start and stop talking at exactly the same times throughout the mixture.
This assumption is easily satisfied: For two speakers to always talk simultaneously, one speaker would have to deliberately start and stop speaking at the exact same times as the other, effectively aiming to sabotage the other’s conversation. Such speakers interfering with intention are unlikely to exist in typical cocktail-party scenario, when interfering speakers in the environment are talking independently from the target speaker. We quantitatively verify this assumption on the real-world audio dataset in Appendix A.
Based on the assumption above, the Positive and Negative Enrollments divide interfering speakers into 4 categories, depending on their existence in both enrollments:
1) Negative Interferer (NI): These are speakers who talk consistently throughout the Positive Enrollment and are present in the Negative Enrollment. The model should learn their voice characteristics from the Negative Enrollment and exclude them from the encoding.
2) Positive Interferer (PI): These are speakers who are not present in the Negative Enrollment and talk in a fraction of the Positive Enrollment. The model should identify those segments when PIs are absent in the Positive Enrollment and exclude these speakers in the encoding. We explore how sensitive our model is to this fractional overlap in Section 4.4.
3) Hybrid Interferer (HI): These speakers appear partially in both the Positive and Negative Enrollments. They resemble a combination of Positive and Negative Interferers, providing the model with two potential cues for suppression: their absence in some Positive Enrollment segments and their presence in some of the Negative Enrollment segments.
4) Neglect-Required Interferer (NRI): These speakers appear exclusively in the Negative Enrollment. Similar to the Negative Interferers, the model encoder should avoid include these speakers’ voice characteristic in the extracted embedding.
Since Positive and Negative Interferers represent the two primary scenarios in which the model must learn to distinguish target speakers from interfering ones, we focus our training and evaluation on these two types. We demonstrate our model’s generalizability when Hybrid and Neglect-Required Interferers exist in Appendix H and I, respectively.
3.3 Model Architecture
Figure 2: Encoding and Extraction Branch model architecture and training pipeline.
As shown in Figure 2, we adopt TF-GridNet tfgridnet as the backbone for both our Encoding and Extraction Branches. The TF-GridNet architecture consists of a 2D convolution layer followed by stacks of three TF-GridNet blocks. Each block consists of two BiLSTM modules that capture the inter-frequency and temporal information, followed by a Full-band Self-attention Module designed to capture the long-range frame information between frames. To adapt the original TF-GridNet, which is designed to perform sound separation, to the Target Speaker Extraction task, we introduce an attention-based Encoder Fusion Module and an Extraction Fusion Module. The Encoder Fusion Module performs comparison between Positive and Negative Enrollment embeddings to generate the embedding of the target speaker, while the Extraction Fusion Module integrates enrollment encoding into the extraction model. The model architecture and training method for encoding the target speaker’s feature are shown in Figure 2.
The Positive and Negative Enrollments are first encoded by a pair of TF-GridNet encoders in the Siamese Encoder. Since The goal of both branches in the Siamese Encoder is to encode multiple speakers’ characteristics from the noisy Positive or Negative Enrollments, we share the parameters between the two TF-GridNet encoders of the Siamese Encoder to reduce model parameter size. The Siamese encoder results in two sequences of embeddings $E_{pos}$ and $E_{neg}$ with shapes $T_{pos},D]$ and $T_{neg},D]$, respectively. $T_{pos},T_{neg}$ are the numbers of frames in the positive and negative enrollment embeddings, and $D$ is the feature dimension at each frame.
# Input:
#   E_pos: shape [T_pos, D]
#   E_neg: shape [T_neg, D]
# Learnable parameters:
#   Segment Embeddings:
#     S_pos, S_neg: shape [1, D]
#   M: 2 Full-band Self-attention layers
Encoder Fusion Module:
  # Step 1: Elementwise add via broadcasting
  E_pos = E_pos + S_pos # shape [T_pos, D]
  E_neg = E_neg + S_neg # shape [T_neg, D]
  # Step 2: Concatenate along the temporal dimension
  E_concat = [E_pos, E_neg] # shape [T_pos + T_neg, D]
  # Step 3: Apply two Full-band Self-attention layers
  E_concat = M(E_concat) # shape [T_pos + T_neg, D]
  # Step 4: Truncate embeddings
  E_fused = E_concat[:T_pos] # shape [T_pos, D]
Figure 3: Encoder Fusion Module pseudo-code.
The Encoder Fusion Module then extracts the target speaker’s voice embedding from the Siamese Encoders’ output $E_{pos}$ and $E_{neg}$. The pseudo-code for the Encoder Fusion Module is shown in Figure 3. In Step 1, two segment embeddings, $S_{pos}$ and $S_{neg}$ with shape $[1,D]$, are first element-wise added to the input embeddings to allow the model to distinguish which enrollment each embedding originates from. The resulting two embeddings are concatenated along the temporal dimension in Step 2, and passed through two Full-band Self-attention layers in Step 3. In these two layers, the self-attention calculation between the embeddings of different Positive Enrollment frames allows the model to identify the Positive Interferers, who remain silent in some of the Positive Enrollment frames, and exclude these speakers in the extracted embedding. Similarly, attention between the Positive and Negative Enrollments embeddings enables the model to identify and exclude Negative Interferers, (i.e., speakers present only in the negative enrollments), from the output embedding. Finally, in Step 4, we truncate the output to retain only the embeddings corresponding to the positive enrollment as the extracted target speaker embedding, which reduces the output to a $T_{\text{pos}},D]$ shaped embedding $E_{\text{fused}}$. This reduction in embedding size ensures the embedding feature is the same size as the pretrained encoder’s output, allowing the model to benefit from the fast convergence brought by training the encoder using knowledge distillation (as shown in Section 4.3). To further reduce the computation load in the extraction branch, we apply non-overlapping average pooling with 40 kernel size along the temporal dimension to $E_{\text{fused}}$, resulting in a $T_{\text{pos}}//40,D]$ shaped feature $E_{\text{fused-pooled}}$. We discuss the effect of different kernel sizes on the model performance and inference time in Appendix D.
Conditioned on the pooled enrollment encoder output $E_{\text{fused-pooled}}$, we use a TF-GridNet-based Extraction Branch to extract the target speaker’s voice from the Audio Mixture. To perform causal target speaker extraction, we modify the TF-GridNet block by adopting the same modification as LookOnceToHear Veluri2024lookonce. Three causal TF-GridNet blocks are used in the Extraction Branch. Two cross-attention-based Extraction Fusion Blocks are added after the first two causal TF-GridNet blocks to integrate the pooled target speaker embedding $E_{\text{fused-pooled}}$ into the extraction model. In each fusion module, the pooled target speaker embeddings $E_{\text{fused-pooled}}$ serve as the Key and Value, while the output from the previous TF-GridNet block serves as the Query for the fusion module. The output from the last TF-GridNet block passes through a transposed convolution layer and an ISTFT module to generate the speech waveform.
In addition to the monaural model described above, which processes single-channel audio, we propose a binaural variant designed to encode and extract binaural audios. The difference between the monaural and binaural model architectures lies solely in the input. The binaural model takes in 4 channels input (stacked real and imaginary components from both channels) and the monaural model takes in only 2 channels (stacked real and imaginary component). By stacking the two channels’ STFT, the binaural model can capture the inter-channel temporal differences between the two channels, which encode the directional information of the target speaker since the target speaker speaks at 90 degree azimuth angle in the enrollment. Both model architectures do not differ after the first convolution layer.
3.4 Training Method
Due to the additional speaker and background noise in the enrollments, the noisy Positive and Negative Enrollment pairs introduce significant variability in the training data in comparison to the clean enrollments. As shown in Section 4.3, when the Encoding and the Extraction branches are trained together using these noisy training samples, the model converges much more slowly and reaches worse final performance. Motivated by the prior works Veluri2024lookonce; splitrole; voicefilter that successfully trained extraction models conditioned on the clean target speaker embeddings, we separately train the Encoding and Extraction branches in a two-stage training strategy.
As shown in Figure 2, the first stage trains the Siamese Encoder and the Encoder Fusion Module from scratch, using knowledge distillation, to extract the target speaker’s embedding from the noisy positive and negative audio enrollments. We obtain the ground-truth target speaker’s embedding by passing the clean target speaker’s voice through a trained frozen TF-GridNet encoder Veluri2024lookonce. Formally, the first stage training loss is written as
$L_{\text{stage 1}}=\|E_{\text{clean}}-E_{\text{fused}}\|^{2}\,$, (2)
where $E_{\text{clean}}$ is the embedding of the clean target speaker voice $a_{\text{clean}}$ encoded by a trained TF-GridNet encoder from Veluri2024lookonce.
In the second stage, we train the Extraction Branch to extract the target speaker from the Audio Mixture. The training loss for the second stage is the negative SNR value of the extracted audio $\hat{a}_{tgt}$ and the ground-truth target speaker speech $a_{tgt}$:
$L_{\text{stage 2}}=-\text{SNR}(\hat{a}_{tgt},a_{tgt}).$ (3)
More training and model details are discussed in Appendix C.
4 Experiment
4.1 Datasets and Baselines
Datasets
We construct our Positive, Negative, and Mixed Audios from the samples in the LibriSpeech dataset librispeech, with background noise $n^{\{M,P,N\}}$ from the WHAM! dataset wham. Following the experiment configuration in LookOnceToHear Veluri2024lookonce, we also construct binaural Positive / Negative / Mixed Audio samples by convolving the speech of each speaker with the binaural RIR data from 4 different datasets: CIPIC cipic, RRBRIR RRBRIR, ASH-Listening-Set ASH-brir, and CATTRIR CATTRIR.
In evaluation, 5000 samples are generated from the LibriSpeech dataset test-clean component using the same strategy as the training samples. We report the average improvement in SNR (SNRi) and SI-SNR (SI-SNRi), and the extracted audio’s STOI and DNSMOS, along with the standard deviation across evaluation samples calculated using the built-in Pytorch function under the assumption that model performance metrics follow a normal distribution. We detail the training and testing data generation in Appendix B.
Baselines
We compare our monaural model’s performance with TCE Chen2024tce, USEF-TFGridnet zeng2024useftse, and non-negative matrix factorization (NMF) method NMF. We compare our binaural model’s reverberant target speech extraction performance with the LookOnceToHear model Veluri2024lookonce.
1. TCE Chen2024tce considers speakers whose voices do not overlap with a given speaker as the target speakers and extracts their voices. The model takes in a d-vector embedding of the given speaker and removes all interfering speakers who talk at the same time as the given speaker.
2. LookOnceToHear Veluri2024lookonce is a binaural target speech extraction method conditioned on noisy audio enrollments. This method uses beamforming to extract the target speaker’s characteristics from the noisy Positive Enrollment, which contains audio from multiple additional speakers.
3. USEF-TFGridnet zeng2024useftse is a target speech extraction model using clean target speaker enrollment as the extraction condition. We include this model as a baseline to show the influence of having interfering speakers in the audio enrollment on the model performance.
4. NMF NMF is a non-deep-learning-based audio separation technique. Since the method could not perform target speaker extraction conditioned on noisy audio examples, we report its extraction quality by selecting the audio with the highest SNR among its output as its extracted audio.
4.2 Result on Monaural and Binaural Reverberant Target Speech Extraction
Table 1: Comparison between our method and the baselines’ performance when extracting from an audio mixture of 2-3 speakers conditioned on an enrollment mixture of 2-4 speakers. NMF’s performance does not change as the number of speakers in the Enrollments changes, as it performs unconditional sound separation. Differences in speaker similarity across combinations substantially impact extraction difficulty, leading to the high standard deviation across samples.
Method Audio Type Metric 2 Speakers in Mixture 3 Speakers in Mixture
2 Speakers 3 Speakers 4 Speakers 2 Speakers 3 Speakers 4 Speakers
in Enroll in Enroll in Enroll in Enroll in Enroll in Enroll
[…]
USEF-TFGridnet zeng2024useftse Mono SNRi 3.42 $\pm$ 3.43 3.45 $\pm$ 3.58 3.30 $\pm$ 3.50 4.31 $\pm$ 2.52 4.15 $\pm$ 2.47 4.23 $\pm$ 2.58
SI-SNRi -0.03 $\pm$ 5.97 -0.03 $\pm$ 6.42 -0.14 $\pm$ 6.00 0.29 $\pm$ 3.38 0.11 $\pm$ 3.36 0.09 $\pm$ 3.24
PESQ 1.52 $\pm$ 0.49 1.54 $\pm$ 0.49 1.52 $\pm$ 0.49 1.32 $\pm$ 0.38 1.32 $\pm$ 0.37 1.31 $\pm$ 0.36
STOI 0.43 $\pm$ 0.17 0.43 $\pm$ 0.18 0.43 $\pm$ 0.17 0.36 $\pm$ 0.11 0.35 $\pm$ 0.11 0.36 $\pm$ 0.11
DNSMOS 1.37 $\pm$ 0.45 1.37 $\pm$ 0.45 1.36 $\pm$ 0.45 1.35 $\pm$ 0.41 1.33 $\pm$ 0.41 1.34 $\pm$ 0.42
WER 0.66 $\pm$ 0.33 0.66 $\pm$ 0.36 0.68 $\pm$ 0.32 0.85 $\pm$ 0.26 0.86 $\pm$ 0.21 0.86 $\pm$ 0.20
[…]
PESQ 1.91 $\pm$ 0.34 1.80 $\pm$ 0.41 1.73 $\pm$ 0.46 1.91 $\pm$ 0.34 1.53 $\pm$ 0.39 1.48 $\pm$ 0.40
[…]
WER 0.73 $\pm$ 0.16 0.76 $\pm$ 0.24 0.77 $\pm$ 0.24 0.73 $\pm$ 0.25 0.88 $\pm$ 0.21 0.88 $\pm$ 0.17
[…]
Ours PESQ 1.84 $\pm$ 0.32 1.83 $\pm$ 0.34 1.84 $\pm$ 0.33 1.57 $\pm$ 0.31 1.56 $\pm$ 0.32 1.54 $\pm$ 0.35
[…]
WER 0.54 $\pm$ 0.37 0.54 $\pm$ 0.29 0.54 $\pm$ 0.28 0.73 $\pm$ 0.24 0.72 $\pm$ 0.25 0.73 $\pm$ 0.24
[…]
Ours PESQ 2.07 $\pm$ 0.34 2.06 $\pm$ 0.36 2.05 $\pm$ 0.36 1.79 $\pm$ 0.33 1.78 $\pm$ 0.34 1.75 $\pm$ 0.37
[…]
WER 0.42 $\pm$ 0.35 0.43 $\pm$ 0.28 0.45 $\pm$ 0.28 0.61 $\pm$ 0.28 0.61 $\pm$ 0.28 0.62 $\pm$ 0.28
We compare our model with baseline methods under scenarios where different numbers of speakers are present. Table 1 shows the monaural target speaker extraction performance. USEF-TFGridnet zeng2024useftse shows significantly worse performance, especially as the number of interfering speaker increase in the enrollment. TCE Chen2024tce achieves better performance than USEF-TFGridnet zeng2024useftse baseline since it does not rely on the clean target speaker’s enrollment. However, apart from the STOI metric in one scenario, the TCE model shows worse performance than our method under all different numbers of interfering speakers in the Audio Mixture and in enrollments.
Table 2: Comparison between our method and the baselines’ performance when extracting binaural audio from mixtures of 2-3 speakers conditioned on an enrollment mixture of 2-4 speakers.
Method Audio Type Metric 2 Speakers in Mixture 3 Speakers in Mixture
2 Speakers 3 Speakers 4 Speakers 2 Speakers 3 Speakers 4 Speakers
in Enroll in Enroll in Enroll in Enroll in Enroll in Enroll
[…]
LookOnceToHear PESQ 2.28 $\pm$ 0.47 2.25 $\pm$ 0.52 2.28 $\pm$ 0.45 1.87 $\pm$ 0.48 1.88 $\pm$ 0.44 1.90 $\pm$ 0.47
[…]
WER 0.45 $\pm$ 0.41 0.44 $\pm$ 0.35 0.47 $\pm$ 0.68 0.66 $\pm$ 0.36 0.63 $\pm$ 0.33 0.64 $\pm$ 0.33
[…]
WER 0.45 $\pm$ 0.48 0.45 $\pm$ 0.34 0.44 $\pm$ 0.35 0.63 $\pm$ 0.33 0.64 $\pm$ 0.43 0.64 $\pm$ 0.42
In the multi-channel target speech extraction, prior work Veluri2024lookonce has explored using beamforming to extract the target speaker’s characteristics. In this experiment, we show that additionally including the Negative Enrollment helps improve model performance. As shown in Table 2, in comparison to the LookOnceToHear Veluri2024lookonce baseline, based on two-sample t-tests over the 5000 test samples at the 90% confidence level, our model achieves statistically significant improvements in SNRi and SI-SNRi under all conditions. The worse STOI performance might be caused by the difference in parameter size, as shown in Appendix C. In contrast to the monaural experiment results, the binaural models show lower performance in most of the metrics and scenarios. This performance drop might be caused by the difference in training objective, since the binaural models need to predict the reverberant binaural audio, while the monaural models are trained to extract non-reverberant speech (to achieve fair comparison with the prior works focusing on non-reverberant extraction). The DNSMOS, PESQ, and WER metric are also affected by the reverberation effect in the prediction, thus providing less effective information for comparing model performance. We include them here only for completeness.
4.3 Ablation Study
Effectiveness of the two-stage model training
Figure 4: Validation loss values for the optimization step. The curve for the two-stage training begins at the 200k step to account for the 200k optimization steps performed during the first training stage.
In this section, we show the effectiveness of the two-stage training by showing the validation loss curve of our model with and without the first training stage. As shown in Figure 4, end-to-end training reaches 3 dB SNR on the validation set after 600k optimization steps (around 125 hours), while the combined training time for the two-staged training takes 240k optimization steps (around 50 hours) based on a single Nvidia A10 24GB GPU.
Such a performance difference might be caused by the high difficulty in encoding the expected target speaker from the noisy enrollment pair. Since its very challenging for the model to draw similarity between the ground-truth single speaker audio and the noisy pair of enrollments input, training using an end-to-end learning method provides little guidance to the encoder on which speaker is expected to be encoded. In contrast, the two-stage training explicitly guides the model to extract the target speaker’s embedding in the first stage, significantly accelerating convergence.
Cross-Attention based Fusion over Film Fusion Method
Film fusion is widely applied in the prior target audio extraction works film-fusion; Chen2024tce; Veluri2024lookonce. However, Film fusion passes the condition embedding through a linear layer, and element-wise multiplies the output with the input tensor to perform fusion. This limits the model to use a fixed embedding size, which might not be capable of encoding the fine-grained details of the target speaker’s voice characteristics. In this experiment, we show that our attention-based fusion method is more optimal for our target speech extraction task.
We keep the rest of the model architecture intact and only change the model fusion method. In particular, we perform global average pooling on the encoder head output along the temporal dimension to obtain an embedding as the condition input to the Film fusion block. As shown in Table 1, our attention-based fusion method achieves higher performance in all application scenarios. This shows that the cross-attention-based fusion method is more preferable for the target speech extraction conditioned on noisy enrollments.
4.4 Model Performance in Challenging Application Scenarios
Model Performance under different PI and NI’s speech length
As the Positive Interferer (PI)’s speech duration in the Positive Enrollment increases, fewer frames are available for the model to distinguish the Target Speaker from the PI. Similarly, reducing the Negative Interferer (NI)’s speech length in the Negative Enrollment results in less available conditional information to remove NI’s voice characteristics in the target speaker encoding. In this section, we explicitly examine how the model performance varies with the duration of the PI and NI’s speech in the enrollment.
We focus on the scenario where one PI and one NI exist in the enrollments, and the Audio Mixture contains the same interferer as those in the enrollments. Let the length of the PI’s speech in the Positive Enrollment be $l_{pos}$, and the NI’s speech in the Negative Enrollment be $l_{neg}$. In Figure 4, we report the average model performance when extracting 2000 test samples with $l_{neg}$ randomly sampled between 1s to 3s and $l_{pos}$ takes different values between 0s to 3s. We evaluate the extracted audio’s SI-SNRi with respect to the target speaker and the PI’s speech in the Audio Mixture. When $l_{pos}=3$, the model fails to distinguish the two speakers as expected, since both speakers have the same speech length in each enrollment. As $l_{pos}$ decreases below 2.7s, the model correctly identifies the target speaker’s characteristic, as shown by the increasing SI-SNRi of the extracted audio w.r.t. the target speaker’s speech and the decreasing SI-SNRi w.r.t. the PI’s speech. Alternatively, we randomly sample $l_{pos}$ between 1s to 3s and evaluate the model performance when $l_{neg}$ takes a value between 0s to 3s. As shown in Figure 5, the model correctly extracts the target speaker and removes the NI’s speech when $l_{neg}$ increases above 0.3s. These experiments demonstrate that the model indeed leverages the differences in speakers’ speech length in both enrollments to encode the target speaker, and could distinguish speakers from as short as 0.3 seconds of misalignment between their speech.
Figure 4: Model performance w.r.t. length of Positive Interferer in the Positive Enrollment.
Figure 5: Model performance w.r.t. length of Negative Interferer in the Negative Enrollment.
Model performance under inaccurate Positive and Negative Labeling
Table 3: SI-SDRi of 10 Sample using Ground-Truth or Inaccurate User Enrollment Labelings.
Enrollment 1 2 3 4 5 6 7 8 9 10
Labeling Source
Ground Truth 10.98 7.60 8.35 11.71 10.31 12.31 11.01 7.03 8.00 11.81
User-Labelings 10.90 7.60 8.40 11.65 10.23 12.14 10.82 7.32 8.04 11.15
$\pm$ 0.04 $\pm$0.05 $\pm$0.02 $\pm$0.04 $\pm$0.07 $\pm$0.12 $\pm$0.24 $\pm$1.32 $\pm$0.10 $\pm$2.10
In the real-world application scenario, inaccurate Positive and Negative Enrollment labeling from the users could lead to unwanted inclusion of the target speaker’s voice in the Negative Enrollments, or including timesteps when the target speaker remains silent in the Positive Enrollment. Both cases violate our model input assumption. In this section, we demonstrate our model’s tolerance to the inaccuracy in the user-provided Enrollments’ labeling.
We randomly select 2-3 speakers and overlay their voices so that each speaker spans 5-7 seconds of a 10-second audio mixture. For 10 such randomly generated samples, we instruct 10 users to label the start and stop timesteps of a certain speaker following a speaker hint, e.g, the same speaker heard in another clean audio example. The segments not labeled by the user are considered negative enrollments. Using the enrollment labeling provided by users, we perform extraction on another audio mixture containing the target speaker. We report the SI-SNRi of each sample across different users’ labeling, as well as the SI-SNRi if the ground truth labeling is used. As shown in Table 3, our model shows little performance degradation when using the inaccurate positive enrollment labeling provided by the user, despite each start and stop timestep of the positive enrollment being on average $0.30\pm 0.19$ seconds away from the ground truth timestep. In addition, the model shows a small variation in metric value across different users’ labeling. This validates our model’s robustness to the inaccuracy in different users’ labeling.
Model Performance on Real-World Audio Mixtures
Table 4: SI-SDRi under Inaccurate User labelings.
Method MOS
Ours (Monaural) 3.35
Ours (FiLM Fusion) 2.60
USEF-TFGridnet 1.45
Unprocessed 2.10
In this section, we compare our model and the baseline methods on naturally recorded real-world audio mixtures. We evaluate the Mean Opinion Score (MOS) of different models using five naturally recorded audio mixtures sourced from Freesound.org and the VoxConverse dataset. These sounds are noisy real-world speech recordings captured in pubs, metro stations, urban areas, and city council meetings. For the evaluation, we manually labelled the Positive and Negative Enrollment and extracted target speech from Audio Mixture clips taken from the same recording. The duration of the enrollment and Audio Mixture clips range from 3 to 8 seconds.
10 participants are asked to rate the original (unprocessed) Audio Mixture, and the audio extracted by three models. Given a textual description of the target speaker (e.g., the female speaker talking over the crowd), they rated the clarity of the target speech relative to background noise on a 1–4 scale. As shown in Table 4, our model achieved a MOS of 3.35, outperforming all baseline methods. These results demonstrate the superiority and practical applicability of our model in real-world target speech extraction scenarios.
5 Conclusion and Limitation
In this paper, we present a method for performing target speech extraction conditioned on noisy enrollments where interfering speakers’ speech overlaps significantly with the target speaker. We optimize our model to encode the target speaker from easily obtained noisy positive and negative enrollments, without using any clean audio enrollment input. Experiments show that our method achieves SOTA performance under the challenging and realistic application scenario.
Though the proposed two-stage training strategy for the encoding branch significantly accelerates convergence, it may limit the encoder’s performance to that of the frozen encoder used for knowledge distillation. Further work on encoding branch training is required to remove such potential performance limitations while maintaining the high convergence speed. In addition, our model extraction still includes artifacts noticeable to the human ear, as shown by the low DNSMOS score reported. This hinders the model’s applicability to downstream tasks such as audio recognition. Further work is required to reduce the artifacts, for example, by employing more advanced extraction branch architectures. Finally, although our model can flexibly perform extraction using clean or noisy audio enrollments, it does not outperform SOTA models explicitly trained to extract conditioned on clean enrollments. Further work is required to develop models with strong generalizability across both enrollment strategies.
The appendix is structured as follow: Section A quantitatively verifies the feasibility of encoding the target speaker using the Positive and Negative Enrollments pair. Section B-D provide training details and justification for the hyperparameter selection. Section E-I present model performance under different numbers and speech lengths of interfering speakers. Section J-L reports the model’s performance when clean audio enrollments are available. Section M show model’s generalizability to speech data from other datasets. Seciton N shows the successful and failure cases. Section O states the societal impact of the work.
Section A :Quantitative Evaluation on the Overlapping Ratio between the Target Speaker and the Interfering Speakers
Section B :Data Simulation Method
Section C :Model Architecture and Training Detail
Section D :Embedding Pooling Size
Section E :Affect of Increasing Number of Positive and Negative Interferers on Model Performance
Section F :Model Performance under Different Lengths of Positive and Negative Enrollment
Section G :Model Performance with respect to the Input Audio Quality
Section H :Model Performance under presence of Hybrid Interferers
Section I :Model Performance under presence of Neglect-Required Interferers
Section J :Model Performance for Target speaker Confusion Problem
Section K :Model Performance when Clean Target Speaker’s Enrollment is available
Section L :Model performance on WSJ0 Dataset
Section M :Audio Visualization and Failure Cases
Section N :Societal Impact
Appendix A Quantitative Evaluation on the Overlapping Ratio between the Target Speaker and the Interfering Speakers
As shown in Section 3.2, we base our problem formulation on the assumption that speakers will not always start and stop talking at the same time. In this section, we show such temporal misalignment between speakers in real-world noisy audio mixtures is prevalent, which allows our model to identify the target speaker from the noisy Positive and Negative Enrollment.
We validate the assumption using the MSDWild dataset and the VoxConverse dataset. Both datasets are activity detection datasets consisting of real-world conversational audios, and provide ground truth activity labeling indicating which speaker speaks when in each conversation. To investigate the overlapping ratio between speakers from different conversations groups, we perform pairwise comparison between conversation samples from each dataset. Specifically, for a pair of conversation containing speakers $\{s_{a1},s_{a2},...,s_{am}\}$ and $\{s_{b1},s_{b2},...,s_{bn}\}$ respectively, we consider all the speaker pairs from different conversation groups $\{(s_{ai},s_{bj})|i\in[1..m],j\in[1..n]\}$, and calculate the mean Intersection Over Union (IOU) for the speech of these speaker pairs $\frac{1}{nm}\sum_{i\in[1..m],j\in[1..n]}IoU(s_{ai},s_{bj})$.
The mean IOU and the standard deviation of IOU across different samples is $0.25\pm 0.20$ for the many-talker val split of the MSDWild dataset, and $0.22\pm 0.06$ for the test split of VoxConverse dataset. This indicates that, on average, the non-overlapping speech between any two independent speakers spans more than half of their combined speech segments. These non-overlapped regions offer rich information for our Positive and Negative Enrollment strategy, enabling the system to identify the target speaker and perform accurate speaker extraction. Note that the non-overlapping regions are defined with respect to pairs of speakers. In scenes with multiple interfering speakers, different interferers may overlap with the target speaker at different times, resulting in a mixture where no single timestep contains only one active speaker. However, by obtaining pairwise speaker difference from the non-overlapping speech between two speakers, Positive and Negative Enrollment pair still provide sufficient information to distinguish the target speaker from interfering speakers.
Appendix B Data Simulation Method
Table 5: Speech length of different types of speakers’ speech in enrollments when 3 seconds of Positive and 3 seconds of Negative Enrollments are used.
Speaker Type Length in Positive Enrollment Length in Negative Enrollment
Target Speaker 3 seconds 0 seconds
Positive Interferer 1-2 seconds 0 seconds
Negative Interferer 3 seconds 1-3 seconds
Hybrid Interferer 1-2 seconds 1-3 seconds
Neglect-Required Interferer 0 seconds 1-3 seconds
In this section, we detail the method for simulating the training and testing audio.
As shown in Section 3.2 in the main paper, based on the presence/absence of the interfering speaker in all/subset of the segments in the enrollments, the interfering speakers are classified into 4 types. Table 5 summarizes the length of each type of speaker in the Enrollments. Target Speaker speaks throughout the Positive Enrollment and does not exist in the Negative Enrollment. Negative Interferers’ speech fully overlap with the target speaker in the Positive Enrollment and exist in the Negative Enrollment. Alternatively, Positive Interferers’ speech does not exist in the Negative Enrollment, and exist in a subset of segments in the Positive Enrollment. Hybrid Interferer exist in a subset of the Positive Enrollment, and also exist in the Negative Enrollment. Neglect-Required Interferers exist only in the Negative Enrollment and are excluded from the Positive Enrollment. In order to ensure each speaker’s voice span the desired length of the Enrollment, we use WebRTC Voice Activity Detector [pywebrtcvad] to detect and remove zeros in the LibriSpeech dataset samples, before repeating or cutting to construct their voice of desired length in the Audio Mixture, the Positive Enrollment and the Negative Enrollment.
We generate our training, validation, and testing data from the train-clean-360, dev-clean, and test-clean components from the LibriSpeech dataset [librispeech], respectively. The sampling rate is 16000. We generate the training samples so that the enrollments and the Audio Mixture all have one target speaker and two interfering speakers. The target speaker is shared between the enrollments and the Audio Mixture, while the interfering speakers in the Audio Mixtures are two randomly sampled speakers different from the target speaker, without being required to be the same as those in the enrollments. To reduce the instability caused by the variation of different speaker types, the Hybrid and Neglect-Required Interferers are not included in the enrollments in the training samples, so the two interfering speakers are randomly designated as either the Positive Interferer or the Negative Interferer. We enforce the target and interfering speakers to talk throughout the Audio Mixture to increase the extraction difficulty. All training samples are generated on the fly to enhance variability and diversity during training. The testing data generation follow the same strategy as the training samples. All the Audio Mixtures in the test samples are 6 seconds long.
To simulate the background environment noise, we use the WHAM! [wham] noise dataset. Different noise samples from the WHAM! noise dataset are recorded in different environments. As a result, additional interfering sounds might exist in one WHAM! noise example but not the other. For instance, some WHAM! noise dataset samples contain music playing in the background while others don’t. Using a WHAM! noise sample containing music in the Positive Enrollment and a WHAM! sample of ambient environment noise in the Negative Enrollment will confuse our model on whether the music is the target sound to be extracted. We resolve such ambiguity by sampling the $n^{\{M,P,N\}}$ from the same WHAM! noise sample but at different temporal segments. Selecting random segments from the WHAM! noise prevents the model from assuming the background noise is the same across the Positive, Negative, and Mixed Audio. We scaled the WHAM! noise to be between [2.5,2.5][-2.5,2.5] SNR with respect to the ground truth target speaker’s voice in each sample.
To simulate the binaural reverberant training and testing samples, we follow the same data simulation method as in the LookOnceToHear [Veluri2024lookonce]. In particular, we convolve each speaker’s voice with BRIR from four binaural room impulse response datasets [cipic, ASH-brir, RRBRIR, CATTRIR]. The binaural RIRs used for constructing one data sample are randomly selected such that 1. All BRIR used in a single data sample generation are from the same scene, and 2. The BRIR for the target speaker in the Positive Enrollment has the direction of arrival of 0 degrees. The pre-processing strategy for each speaker’s voice and the generation of background noise remains the same as the monaural dataset.
Appendix C Model Architecture and Training Detail
Table 6: Comparison of model parameter size, inference time, and memory usage when extracting 1-second audio on an Intel Xeon Silver 4314 @ 2.40GHz CPU. We report the average inference time for extraction in 50 repeated experiments. Replacing the Film fusion module with the cross-attention based fusion module in our final model architecture could significantly reduce the parameter size.
Model Param Size Inference Time Inference Memory Usage
Ours (cross attn) 1.88 M 0.36s 2.30 GB
Ours (Film fusion) 3.94 M 0.3s 2.31 GB
TCE 2.54 M 0.11s 1.89 GB
LookOnceToHear 4.41 M 0.35s 2.32 GB
The TF-GridNet [tfgridnet] Encoder in the Encoding Branch uses the following configuration: $4\times 4$ kernel size and $1\times 1$ stride in the first Conv2D, 64 hidden units in all three BiLSTM layers, 8 attention head numbers in the Full-band Self-attention Module. The input audio is processed by Short-Time Fourier Transform (STFT) with 128 window size and 64 hop length.
To perform causal inference in the extraction branch, we follow the same modification as in the LookOnceToHear [Veluri2024lookonce]. In particular, we remove the global layer normalization after the first convolution layer, change the BiLSTM in the TF-GridNet blocks to unidirectional LSTM, and constrain the Full-band Self-attention Modules to calculate the causal attention value of a one-time frame with only frames before it. The parameter configurations of causal TF-GridNet blocks in the Extraction Branch are the same as the Encoding Branch, apart from using $1\times 1$ kernel size in its first Conv2D layer. Three causal TF-GridNet blocks are used in the Extraction Branch. The output from the last TF-GridNet block passes through a transposed convolution layer and an ISTFT module to generate the speech waveform.
All the training is done on a single Nvidia A10 24GB GPU with a batch size of 2. In both training stages, we use the Adam optimizer and decay the learning rate by half when the validation loss does not decrease for more than 50 epochs. Motivated by MOCOv3 [mocov3], which observes that changes in the parameters of shallower layers can lead to instability in the loss, we assign lower learning rates to the earlier layers of the model and higher learning rates to the deeper layers. In particular, the initial learning rate is set to 5e-4 for the whole Siamese Encoder, 1e-3 for the Encoder Fusion Module in the first training stage, and 2e-3 for the whole extraction branch in the second training stage. 500 epochs (200k optimization steps) are used in the first pretraining stage, and 1000 epochs (400k optimization steps) are used in the second stage.
Since TCE is not trained on the LibriSpeech dataset, we fine-tune the TCE model on our simulated data till convergence (for 120k optimization steps) using the Adam optimizer with 5e-4 learning rate. To train and test the TCE model on the proposed extraction task, we modify our dataloader’s output to match TCE’s application scenario by adding a known speaker’s voice in the Negative Enrollment and concatenating the Mixed Audio with this modified Negative Enrollment. The model then performs extraction conditioned on the known speaker’s d-vector embedding.
In addition, despite the LookOnceToHear model is trained on audio examples from the LibriSpeech dataset, it uses Scaper toolkit to load and generate audio mixtures. We notice this leads to slightly different audio input to the model if the audio is otherwise loaded by torchaudio.load and mixed by addition. The distribution shift in input audio leads to a noticable difference of around 1 SI-SNRi decrease in performance. For fair comparison, we tune the LookOnceToHear’s extraction model on our training data with an initial learning rate set to 5e-4 and decay by half if the validation error does not decrease for more than 50 epochs. The model performance converges after 300 epochs fine-tuning (120k optimization steps) when the lr decreased to around 6e-5.
We load the USEF-TFGridnet model from the open-source project https://github.com/ZBang/USEF-TSE.
In Table 6, we compare our model architecture’s parameter size, extraction branch’s inference time, and memory usage in inference, with the two prior works on target speaker extraction using noisy enrollments. Our model uses smaller parameter size. In particular, we notice that the Film fusion module substantially increases the parameter size, as shown by the significant increase in model parameter size when we replace our cross-attention fusion module with the Film fusion module.
Appendix D Embedding Pooling Size
Table 7: Model performance with different pooling sizes. The inference time is the average time taken for the model main branch to extract 1-second audio in 50 repeated experiments on an Intel Xeon Silver 4314 @ 2.40GHz CPU.
Pooling Size Metric 2 Speakers in Mixture 3 Speakers in Mixture Inf.
2 Speakers 3 Speakers 4 Speakers 2 Speakers 3 Speakers 4 Speakers Time
in Enroll in Enroll in Enroll in Enroll in Enroll in Enroll
[…]
To reduce the model computation time in the Extraction Branch, we perform an average pooling of 40 on the embedding extracted by the encoder head. In this section, we analyze the trade-offs between model performance and inference time under different pooling configurations. Specifically, we compare our model (40 pooling size) with two variants using pooling sizes of 20 and 80, as shown in Table 7.
As the pooling size increases, the length of the extracted target speaker embedding sequence decreases, which reduces the inference time for the extraction branch. Larger pooling size also results in more information loss, which leads to worse extraction performance. To our surprise, a smaller pooling size also results in a worse performance, despite it retains more detailed information of the target speaker in the embedding sequence. We hypothesize that smaller pooling size introduces instability in model training, as the resulting target speaker embedding sequence has larger variation across different timesteps. The unstable model learning could lead to worse final model performance. As a result, we selected 40 pooling sizes as the final model configuration.
Appendix E Affect of Increasing Number of Positive and Negative Interferers on Model Performance
Table 8: Model performance under different number of Positive Interferers (PI).
Method Metric 2 Speakers in Mixture 3 Speakers in Mixture
1 PI 2 PI 3 PI 1 PI 2 PI 3 PI
[…]
Table 9: Model performance under different number of Negative Interferers (NI).
Method Metric 2 Speakers in Mixture 3 Speakers in Mixture
1 NI 2 NI 3 NI 1 NI 2 NI 3 NI
[…]
In Section 4.2 in the main paper, we randomly assign the interferers in the enrollments as Positive or Negative Interferers, and show the model performance as the number of interferers increase. In this section, we separatedly evaluate how Positive and Negative Interferers will affect model performance. In the first experiment, we include only Positive Interferers in the enrollments and gradually increase their number. In the second experiment, we follow the same procedure but only include Negative Interferers in enrollments. In both cases, we evaluate the model’s performance when extracting a target speaker from Audio Mixtures containing 2–3 speakers.
As shown in Table 8 and Table 9, our model shows only 0.2 dB SI-SNRi decrease in performance as the number of Positive Interferer increases from 1 to 3. In comparison, when the number of Negative Interferers increase, the model shows more significant performance decrease of over 1 dB SI-SNRi. This suggests that, in comparison to removing interfering speakers’ characteristic given in the Negative Enrollment, our model is more capable of identifying the target speaker’s characteristic, which is a common voice characteristic that exists across different segments in the Positive Enrollment.
Appendix F Model Performance under Different Lengths of Positive and Negative Enrollment
Table 10: Model performance under different monaural enrollment lengths. We vary the conditional Positive and Negative Enrollment length between 1 to 10 seconds.
Condition Metric Pos. 1 sec Pos. 3 sec Pos. 5 sec Pos. 10 sec
[…]
We train our model solely on samples with 3 seconds of Positive and Negative Enrollments, and tested the model performance solely on these scenarios. In this section, we test the model performance on different enrollment lengths between 1 to 10 seconds. We keep the ratio of the Positive Interferer in the Positive Enrollment, and the ratio of the Negative Interferer in the Negative Enrollment unchanged. This means the Positive Enrollment will still span $1/3$ to $2/3$ of the Positive Enrollment, and Negative Interferer span the $1/3$ to full length of the Negative Enrollment. As shown in Table 10, the model performance improves in general as both the length of the Positive Enrollment and the Negative Enrollments increase. We also notice that increasing the Negative Enrollment length when less than 3 seconds of the Positive Enrollment is provided could adversely affect the model performance. This might be because the model has too little information of the target speaker form the Positive Enrollment, while the Negative Enrollment might guides the model to remove excessive speaker characteristics from the Positive Enrollment, resulting in the decrease in model performance. This suggests that users can enhance extraction quality by providing longer Positive and Negative enrollments if the results are unsatisfactory, and increasing the Positive Enrollment length is preferable.
Appendix G Model Performance with respect to the Input Audio Quality
Figure 6: SI-SNR of the extracted audio with respect to the Mixed Audio. The dashed black line represents the zero-improvement line.
Figure 7: SI-SNRi of the extracted audio with respect to the Positive Enrollment SI-SNR.
Model Performance with respect to the Audio Mixture Quality
In this section, we show our model performance relative to the quality of input Audio Mixture. As shown in Figure 7, we plot the extracted audio’s SI-SDR values for 5000 randomly generated test samples as a function of their input SI-SDR. 96.6% of the audio mixtures’ SI-SNR were improved after the extraction.
Model Performance with respect to the Positive Enrollment Quality
Figure 7 presents our model’s extraction performance as a function of enrollment audio quality. To evaluate this, we scale the clean enrollment audio of the target speaker (Positive Enrollment) to specific SI-SNR values, while leaving the Negative Enrollment and the Mixed Audio unchanged. We then measure the average of 5000 test samples’ extracted audios’ SI-SNRi at each Positive Enrollment SI-SNR value.
The model achieves optimal performance when the Positive Enrollment SI-SNR is between -5 dB to 5 dB, and the performance deteriorates significantly when the SI-SNR of the Positive Enrollment falls below -10 dB. To address this limitation, we fine-tune our model by scaling the training data’s Positive Enrollment to -20 and -15 SI-SNR dB. As shown by the red line and the orange line in Figure 7, this fine-tuning shifts the model’s peak performance to lower SI-SNR regions, which means the model achieves optimum extraction performance under worse Positive Enrollment quality. Further research is necessary to develop a model with strong performance across the full SI-SNR range of the Positive Enrollment.
Appendix H Model Performance under presence of Hybrid Interferers
Table 11: Model performance under different number of Hybrid Interferers (HI)
Method Metric 2 Speakers in Mixture 3 Speakers in Mixture
1 HI 2 HI 3 HI 1 HI 2 HI 3 HI
[…]
As defined in Section 3.2 in the main paper, when an interfering speaker speaks in a subset of segments in the Positive Enrollment, and also exist in the Negative Enrollment, it is a Hybrid Interferer. Model have two ways to distinguish these interfering speakers from the target speaker: either through noticing that these speakers were silent in some of the segments in the Positive Enrollment, when the target speaker is supposed to talk continuously; or noticing that these speakers exist in the Negative Enrollment, when the target speaker is supposed to remain silent. In this section, we explicitly show the model performance when such interfering speakers exist.
As shown in Table 11, we focus on the scenario containing only Hybrid Interferers, and gradually increase their number from 1 to 3. In comparison to the increase in the Positive and Negative Interferer number, our model shows higher robustness to the increase in the number of Hybrid Interferers. This indicates that our model generalizes well in removing interference from Hybrid Interferers by leveraging the fact that such speakers can be effectively suppressed when treated as either Positive or Negative Interferers.
Appendix I Model Performance under presence of Neglect-Required Interferers
Table 12: Model performance under different number of Neglect-Required Interferers (NRI).
Method Metric 2 Speakers in Mixture 3 Speakers in Mixture
1 NRI 2 NRI 3 NRI 1 NRI 2 NRI 3 NRI
[…]
When recording the Negative Enrollments, additional interfering speakers not captured in the Positive Enrollments might be present. Since these interferers’ speech does not overlap with the target speaker, their voice should be neglected (thus named neglect-Required Interferers). In this section, we investigate their affect on our model performance in this section.
As shown in Table 12, similar to the result observed when the number of Hybrid Interferers increase, our model does not show statistically significant performance differences when the number of Neglect-Required Interferers increases from one to three. This demonstrates our model effectively generalizes to prevent the Neglect-Required Interferers from adversely affecting the quality of the extracted target speaker speech.
Appendix J Model Performance when Clean Target Speaker’s Enrollment is available
Table 13: Comparison between our method and the TD-SpeakerBeam baselines’ performance when extracting conditioned on clean or noisy enrollments.
Method Metric 2 Speakers in Mixture 3 Speakers in Mixture 4 Speakers in Mixture
Clean Enroll Noisy Enroll Clean Enroll Noisy Enroll Clean Enroll Noisy Enroll
[…]
If applied without modification, our model fails when the enrollment consists solely of the clean target speaker’s voice. This is due to the model’s input normalization, where an all-zero Negative Enrollment leads to a NaN output. To address this, we randomly select a fixed WHAM! noise sample as the Negative Enrollment, and add it with the Positive Enrollment to construct a pseudo-noisy Positive Enrollment. Using such modification, we construct pseudo-Positive and Negative Enrollment from different target speakers’ clean audio enrollment, and perform extraction.
As shown in Table 13, in comparison to extracting target speaker’s characteristic from noisy Enrollments, our model achieves higher performance when extracting conditioned on the pseudo-Positive and Negative Enrollment. In comparison to the TD-SpeakerBeam [speakerbeam], which is trained to extract using clean audio enrollment, our model achieves higher SNRi, SI-SNRi, and STOI when extracting from the mixture of 3 or more mixtures. These results verifies our model’s applicability to extraction using clean audio enrollments. The reason why our model achieves higher performance under larger number of interfering speaker might be the difference in training data generation. TD-SpeakerBeam is trained on predefined mixtures of different speakers’ speech, while our training samples are generated by dynamic mixing, which creates more diverse mixture of different speakers’ voices. This training data generation strategy acts as a form of data augmentation, helping the model generalize better to larger number of interfering speaker. However, there is still noticeable performance gap between our model and the TD-SpeakerBeam when extracting from the mixture of two speakers using clean enrollments. Further work is required to reduce the performance gap between models using noisy enrollment and using clean enrollment.
We want to emphasize that our model tackles the more challenging task of target speech extraction (TSE) with noisy enrollment, a realistic scenario where prior works have shown significantly degraded performance. Although our model has worse performance under clean audio enrollment, our method still correctly identifies the target speaker users want, shown by the strongly positive SI-SNRi values. Indeed, the performance gap between our method and other models under clean enrollment could bring novel insights to improve our model, which we will explore in the further works.
Table 14: Model performance when extracting conditioned on noisy single-speaker enrollments.
Method Metric 2 Speakers 3 Speakers
in Mixture in Mixture
[…]
Table 15: Speaker configuration for evaluating the Target Speaker Confusion Problem.
Target Speaker Interfering Speaker
Audio Mixture A B
$P_{pred}$ A B
$P_{tgt}$ A B
$P_{interfer}$ B A
In addition, we evaluate our model and the baseline models’ performance when using noisy single speaker enrollments. The noise in the enrollment is from the WHAM! dataset and set at 0 SNR level. As shown in Table 15, TSE models trained with clean enrollments show significant performance decrease under noisy single speaker enrollments. In comparison, our model consistently outperforms the baselines across all metrics except DNSMOS, demonstrating greater robustness to enrollment noise.
Appendix K Model Performance for Target speaker Confusion Problem
Prior works have shown that the auxiliary speaker encoder may sometimes generate ambiguous speaker embeddings, leading to the extraction model extracting interferer’s voice as the output. This is known as the target confusion problem [zhao2022targetconfusion]. To further investigate whether our model suffers from the target speaker confusion problem, we follow the experimental setup proposed by Zhao et al. [zhao2022targetconfusion]. Specifically, we construct 5000 test samples where both the audio mixture and the enrollment contain two speakers, and share the same interfering speaker. Let $P_{pred}$ be the enrollment pair in the sample, where speaker A is the target speaker and speaker B is the interfering speaker. After performing extraction, we identify the enrollment pairs that result in the extracted audio being more similar (in terms of SNR) to the interfering speaker B’s speech than to the target speaker A’s. We refer to these samples as target confusion samples. For each target confusion sample, we construct two additional enrollment pairs: $P_{tgt}$ and $P_{interfer}$. In $P_{tgt}$, same as the target confusion sample, speaker A is the target and speaker B is the interferer. In $P_{interfer}$, speaker B is the target and speaker A is the interferer. Table 15 below summarizes the role of speaker A and speaker B in the audio mixture and each enrollment pair.
To verify if the model mistakenly encode the interfering speaker as the target speaker and the interfering speaker, we flatten and normalize the embeddings extracted by our model from each enrollment pair, obtaining $E(P_{pred})$, $E(P_{tgt})$, $E(P_{interfer})$, and compute the cosine similarity between $E(P_{pred})$ and $E(P_{tgt})$, as well as between $E(P_{pred})$ and $E(P_{interfer})$.
Out of the 5000 tested samples, 74 of them are the target confusion samples. 34 samples out of the 74 target confusion samples (45.9%) have enrollment embedding closer to the interfering speaker than to the target speaker (i.e. 34 of the samples have $cos(E(P_{pred}),E(P_{tgt}))<cos(E(P_{pred}),E(P_{interfer}))$). These findings are consistent with those reported in the target speaker confusion study [zhao2022targetconfusion], where 45.1% of target confusion samples showed the extracted embedding being closer to the interferer. This means that significant percentage of target confusion samples have encoder embeddings being more close to the interfering speaker. However, it is important to emphasis that only 74 out of the 5000 samples (1.48%) showed evidence of target speaker confusion in our evaluation, suggesting that this is a rare occurrence and does not pose a significant issue for our model’s overall performance.
Appendix L Model performance on WSJ0 Dataset
Table 16: Performance of our model and the TCE model before and after fine-tuning, using monaural samples generated from the WSJ0 dataset.
Method Metric 2 Speakers in Mixture 3 Speakers in Mixture
1 NI 2 NI 3 NI 1 NI 2 NI 3 NI
[…]
In the main paper, we focus our experiments on the samples constructed from the LibriSpeech dataset. In this section, we show our model’s applicability to other datasets. We select the WSJ0 dataset as the source of speaker speech.
As shown in Table 16, our model correctly encode and extract the target speaker’s voice, as shown by the positive SNRi and SI-SNRi metric, and outperforms the TCE baseline in most scenarios. However, after fine-tuning on samples containing two interfering speakers in both the Audio Mixture and enrollments, our model only improves performance when extracting from three-speaker mixtures, while performance degrades in two-speaker cases. Further exploration is required to train a model with strong generalizability across multiple datasets and scenarios.
Appendix M Audio Visualization and Failure Cases
Figure 8: Successful cases of our model. We selected four samples where the model SI-SNRi performance is above 11 dB. We report the extracted audio SI-SNRi under the case number.
Figure 9: Failure cases of our model. We selected four samples where the model SI-SNRi performance is below 8 dB. We highlight the timesteps where the model has poor performance with red boxes.
In this section, we show four successful cases and four failing cases of our model. For each case, we visualize the waveform and STFT spectrogram of the Mixed Audio, the extraction ground truth, and the model prediction of a one-second audio segment.
As shown in Figure 8, our model correctly predicts silent sound in the extracted audio when the target speaker is not speaking and extracts the target speaker’s voice despite the frequency range largely overlaps with the interfering speakers.
In several cases, our model has less optimum performance, with less than 7 dB SI-SNRi performance. We show four of these cases in Figure 9. In these cases, the model fails to extract the full target speaker’s voice in some timesteps (Failure Cases 2), includes interfering the speaker’s voice in its extraction (Failure Case 1, 3, 4), and introduces artifacts that obscure the intelligibility of the target speech. These non-exhaustive failure cases serve as a reference for future improvements.
Appendix N Societal Impact
This paper presents work whose goal is to advance the field of Machine Learning and Audio Processing. This technology can potentially improve communication aids, hearing devices, and audio editing tools, making it easier for users to focus on desired voices in complex auditory scenes. We acknowledge that such technology could raise privacy concerns if misused to extract a speaker’s voice without consent. However, our method requires prior access to the target speaker’s speech activity labels, which presupposes that the target speaker is already engaged in a face-to-face conversation or otherwise clearly audible to the user. As a result, our approach serves to enhance human auditory perception rather than introduce new capabilities that could pose societal or ethical risks. While our work may have broader societal implications, we do not identify any that require specific emphasis.